The length of the base of a triangle is twice its height. If the area of the triangle is 196 square​ kilometers, find the height.
Answer by Boreal(15235)

A=(1/2)bh
196=(1/2)bh
392=bh
b=2h
392=2h^2
196=h^2
h=14 km
b=28 km
their product is 392 km^2, and half of that is 196 km^2.